OpenStudy (anonymous):

i cant plot 2295 and (325)²/16sinθcosθ on a graphing calculator!! help me!!

OpenStudy (anonymous):

Here, be sure your calculator is in radian mode. It's like the last one we did. Except here y0=2295 (It's just a constant). y2=325^2 / (16sinxcosx)

OpenStudy (anonymous):

okay i typed that but when i clicked graph, it didnt show up! i dont know why...

OpenStudy (anonymous):

Because those numbers are very very very big. Your graph (by default) goes from -10 to 10 on both axes. When you're on the graph, hit the "zoom" button. Then arrow down to "zoom fit". (It will take a few seconds)

OpenStudy (anonymous):

Once you see it, you'll see that the numbers are extremely large. Don't ask for intersection points since there are an infinite number of them ;]

OpenStudy (anonymous):

okay! its really slow...

OpenStudy (anonymous):

It is! It's working with very very large numbers and it's not a particularly powerful computer :P

OpenStudy (anonymous):

It shouldn't take more than 30 seconds. If it does, turn it off and turn it back on. Sometimes they freeze up.

OpenStudy (anonymous):

because i was clicking a lot of things i got a graph that looks like this now....

OpenStudy (anonymous):

Oh my. Well, that's not correct. Be sure the calculator is in radian mode. y0 = 2295, y1 = 325^2 / 16sinxcosx. Hit graph. Once you see the graph, hit the zoom button, then arrow down to zoom fit.

OpenStudy (anonymous):

it is in radian mode!!

OpenStudy (anonymous):

 What it should look like.